SourMunchkin7806:

Find a cubic function with the given zeros. -2, 5, -6

SourMunchkin7806:

f(x) = x3 + 3x2 + 28x - 60 f(x) = x3 - 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x + 60

SourMunchkin7806:

State the vertical asymptote of the rational function. f(x) = quantity x minus four times quantity x plus nine divided by quantity x squared minus one x = 1, x = -1 None x = 4, x = -9 x = -4, x = 9

iosangel:

(x +2) (x - 5) ( x + 6) [ x^2 + 2x - 5x - 10 ] [ x + 6 ] [ x^2 - 3x - 10] [ x + 6 ] x [ x^2 - 3x - 10 ] + 6 [ x^2 - 3x - 10 ] x^3 - 3x^2 - 10 x + 6x^2 - 18x - 60 combine like terms x^3 + 3x^2 - 28x - 60

SourMunchkin7806:

awesome thanks can you help me with 3 more?

iosangel:

I'll try....

SourMunchkin7806:

State the vertical asymptote of the rational function. f(x) = quantity x minus four times quantity x plus nine divided by quantity x squared minus one x = 1, x = -1 None x = 4, x = -9 x = -4, x = 9

SourMunchkin7806:

i cant find this one

eatsalad:

\[f(x)=\frac{ (x−4)(x+9) }{ (x^2−1) }\] To find vertical asymptote, set the denominator to zero then find 'x' \[x^2-1=0\] \[x^2=1\] \[x=1,x=-1\] Answer: \[x=1,x=−1\]